OpenStudy (anonymous):

20x+15+20(x-1)+15(x-1)=245

OpenStudy (anonymous):

20x+15+20x-20+15x-15-245=0 55x-265=0 55x=265 x=265/55 x=53/11 Hope this helps! :)